OpenStudy (anonymous):

find the equation of the tangent line to the graph of f(x)=cot(2x) at (π/4, 0)

hero (hero):

Use Tangent line approximation for this

hero (hero):

\[y = f(c) + f'(c)(x-c)\]

hero (hero):

\[y = f(\frac{\pi}{4}) + f'(\frac{\pi}{4})(x-\frac{\pi}{4})\]

hero (hero):

\[y = -2x + \frac{\pi}{2}\]

OpenStudy (anonymous):

i think my teacher wants me to find the derivative of cot(2x) first

OpenStudy (anonymous):

yes hero did that as well, just didn't write it

hero (hero):

\[f'(x) = \frac{-2}{\sin^2(2x)}\]

hero (hero):

I fogot

OpenStudy (anonymous):

how did u find the derivative?

OpenStudy (anonymous):

@shortie is it clear what you need to do here? you want an equation for a line. to find an equation of a line you need two things 1) a point 2) the slope you have the point, it is \[(\frac{\pi}{4},0)\] so now you need the slope

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

to find the slope you need to a) take the derivative, which gives you a formula for the slope b) plug in the x - value so that you actually get the number for the slope

OpenStudy (anonymous):

then you use the point - slope formula or its equivalent, the formula that hero wrote up top

OpenStudy (anonymous):

\[y = f(c) + f'(c)(x-c)\] is the point - slope formula rewritten. the point is \[(c,f(c))\] and the slope is \[f'(c)\] note that c and f(c) are numbers.

OpenStudy (anonymous):

Good job hero

OpenStudy (anonymous):

so when I plug (π/4) in, I'm going to get sin²(π/2) in the denominator?

OpenStudy (anonymous):

i guess to but that is lucky because it is just 1, and so your slope is -1

OpenStudy (anonymous):

no that is wrong, your slope is -2, sorry

OpenStudy (anonymous):

hero wrote the answer above. you see that you have a line with slope -2. found by taking the derivative and plugging in the x - value

OpenStudy (anonymous):

I got it! Thanks :)